# pnasresearcharticle

### v1

@@ A Measure of Departure Time Peakedness
We apply the proposed measurement framework to the departure time data from a 12-month (July 2015 to June 2016) record of transit smartcard, encompassing 5,947,907 bus journeys made by 29,640 individuals across three study networks: Brisbane, Gold Coast, and Sunshine Coast (see SI Section 2). We first examine and compare the overall system-wide departure time patterns and the peakedness characteristics of these networks. Fig.1 shows the system-wide departure time distribution for each network, combining all bus commuters’ journeys over a one-year period. The departure time distribution is depicted as a normalised histogram with a bin width of 5 minutes. In Brisbane, two distinct peaks (AM and PM) are observed. Gold Coast and Sunshine Coast display similar dual-peak patterns, although the PM peaks are less sharp compared to Brisbane. Given these two distinct peaks, we separately analyse the peakedness characteristics of the AM and PM periods. Trips occurring between 12:00 AM and 11:59 AM are classified as AM trips, while trips between 12:00 PM and 11:59 PM are classified as PM trips. For each time-of-day group, we estimate three peakedness measures—$\overline{\psi_{h}}$, $\psi_{h}^{sys}$, and $PCF_{h}$—across the three study networks. Fig.4 shows these estimates for different peak window widths, $h\in\{5,10,20,30,45,60\}$. Across the cities, Brisbane exhibits overall higher $\overline{\psi_{h}}$ and higher $PCF_{h}$, and thus higher $\psi_{h}^{sys}$, than Gold Coast and Sunshine Coast in both the AM and PM periods. This is likely due to Brisbane’s higher population density and greater urbanisation, which create more structured, work-driven commuting patterns, resulting in more synchronised and concentrated departure times. On the other hand, the more leisure-oriented nature of Gold Coast and Sunshine Coast tends to result in more flexible and dispersed departure patterns. Between the AM and PM periods, peakedness measures are overall higher during the AM period than during the PM period in each city. This can be attributed to the stricter time constraints in the morning, with commuters needing to adhere to fixed start times for work or school, leading to a greater concentration of AM departure time distributions around their peaks. However, Brisbane’s $PCF_{h}$ measures display an interesting variation: $PCF_{h}$ of PM trips is higher than or similar to that of AM trips, indicating that users’ peak departure time windows coincide more strongly in the PM period than in the AM period, while the other two cities show much lower peak coincidence in the PM period. This can be explained by a more uniform end-of-day schedule across work sectors and after-school activities in Brisbane, resulting in greater synchronisation of departure times across individual users.
@@
Choice of Peak Window, $h$
Selecting the appropriate width of the peak window, $h$, is important in measuring and interpreting $\psi_{h}$, especially when comparing its values across individual users. If $h$ is too large, $\psi_{h}$ will tend to approach 1, as most departures fall within the wide window, leading to an inflated peakedness for most travellers. Conversely, if $h$ is too small, $\psi_{h}$ will approach 0, since very few departures times fall within the narrow window, resulting in artificially low peakedness for most travellers. A desirable property of $h$ is that the resulting $\psi_{h}$ effectively differentiate between different users’ varying departure time patterns. As such, we propose to determine the optimal window $h^{*}$ as the $h$ value that maximises the variance of $\psi_{h}$ ($\text{Var}[\psi_{h}]$) across users (see SI Sections 1.E-1.F. The $h$-min Peak Trip Concentration under this optimal $h^{*}$ ($\psi_{h^{*}}$) can be considered the most discriminating of all $\psi_{h}$ in that it detects the most diverse range of departure time peakedness in the system, thereby allowing us to recognise and distinguish different departure time distributions of different users more effectively.
@@
Analysis for Brisbane, Gold Coast, and Sunshine Coast reveals that the optimal window $h^{*}$ is consistently 20 minutes, regardless of the time-of-day group or region (see SI Section 1.E). This suggests that there is a common temporal threshold where departure time peakedness is most distinguishable across the population. As such, we consider $h$=20 minutes as the most effective window to measure trip concentrations ($\psi_{h}$), as it reveals the greatest diversity in departure time peakedness. Table 1 presents the descriptive statistics of the distribution of estimated $\psi_{h}$ values with $h$=20 minutes (i.e., $\psi_{20}$). Across the three cities, on average, approximately 61-63% of daily departure times occur during each user’s most concentrated 20-minute window in the AM period, and 46-49% in the PM period. Users in the top 5% for peakedness, represented by the 95th percentile, show that nearly 98-100% of their departures are concentrated within the peak 20-minute window in the AM and 92-97% in the PM. In contrast, users in the bottom 5% for peakedness, as indicated by the 5th percentile, have only 20-23% of their departure times within the peak 20-minute window for the AM period, and 16-17% for the PM period.
@@
Factors Affecting Departure Time Peakedness
@@
First, we explored how different passenger types shape departure time peakedness. Fig.5A compares the three peakedness measures across six passenger types: Adult, Tertiary Student, Pensioner, Secondary Student (15+ years), Child (5 to 14 years), and Senior. Notably, the child passenger type exhibits consistently high departure time peakedness across all three measures, particularly in user-level peakedness ($\overline{\psi_{h}}$). For example, in Brisbane, the $\overline{\psi_{h}}$ of children is nearly 0.8, indicating that most of their day-to-day departures occur within their most frequent 20-minute departure time window. This high concentration is likely due to the fixed start and end times of primary and secondary school, which children in the 5-14 year age group typically attend. Interestingly, the peak window coincidence ($PCF_{h}$) for the PM period is notably high for children in Brisbane ($PCF_{h}$=0.6) and Sunshine Coast ($PCF_{h}$=0.5), even surpassing the AM period. This could be attributed to the more uniform scheduling of after-school activities or programs across various areas in these cities, which may lead to a higher alignment of peak departure times among different users within the peak window. This finding provides some evidence to the role of deeply ingrained travel habits in shaping the travel choice of children (29) alongside other factors including service availability, perceived safety and accessibility.
Next, Fig.5B draws together a suite of situational variables known to impact bus ridership dynamics in the form of six scenarios: Weekday vs Weekend, Non School Holiday vs. School Holiday, and Rain vs. No Rain. We find that the departure time peakedness characteristics are relatively stable across these scenarios, with some subtle shifts. For example, in Brisbane and Gold Coast, weekend departure times are slightly less peaked than weekdays, likely due to more flexible weekend schedules with diverse non-work activities, leading to more dispersed travel times. In contrast, weekdays are dominated by structured work and school schedules, resulting in higher peakedness. Sunshine Coast does not show such distinction, possibly due to its tourism-driven, leisure-oriented travel patterns and flexible local employment, resulting in consistent travel behaviour across both weekdays and weekends.
@@
Figure 5: Comparison of peakedness measures ($h$ = 20 minutes) under different passenger types and scenarios
@@
Temporal Stability and Variability in Departure Time Peakedness: A Trend Analysis
@@
Next we investigate long-term temporal dynamics of departure time peakedness to understand whether system- or user-level peakedness characteristics change over time. Analysis focused on tracking trends over a one-year period, from July 2015 to June 2016, using three-month moving windows to create time series of both system-level ($\psi_{h}^{sys}$) and user-level ($\psi_{h}$) peakedness measures. To evaluate the presence of trends, we applied the Mann-Kendall (MK) Trend Test (30, 31), which assesses whether there is a statistically significant monotonic increase or decrease in the peakedness measures over time (see SI Section 3 for detailed methods and results). At the system level, results show no significant monotonic trends in any city (Brisbane, Gold Coast, or Sunshine Coast), indicating that system-level departure time peakedness remains stable over time. This suggests that the collective travel behaviour of long-term regular users does not exhibit major fluctuations in departure time patterns throughout the year. However, at the user level, the analysis reveals notable dynamics. Approximately 25% of long-term regular users (those making trips consistently throughout the year) exhibit significant trends in their $\psi_{h}$ time-series. Interestingly, these trends are balanced: about half of the users show increasing peakedness (with more concentrated and regular departure times over time), while the other half display decreasing peakedness (with more dispersed and flexible travel patterns over time). Some users exhibit gradual changes, while others experience sudden shifts. Overall, this trend analysis highlights the stability of system-level peakedness, while revealing the evolving nature of individual-level departure time behaviour. It also suggests that habitual travel patterns, though resistant to short-term situational factors like weather or holidays, may shift over the long term due to larger-scale factors such as seasonality, service changes, or shifts in user characteristics.
Importance of Distinguishing System-level and User-level Departure Time Peakedness
@@
User-level peakedness and system-level peakedness do not always exhibit a direct correlation. In some cases, areas with users whose departure time distributions have a high degree of peakedness may still result in a combined departure time distribution at the system level that shows low peakedness. This discrepancy arises because the system-level distribution peakedness ($\psi_{h}^{sys}$) is influenced by both the peakedness of individual departure time distributions ($\overline{\psi_{h}}$) and the alignment of their peak time windows ($PCF_{h}$). In situations where the peak windows of individuals are widely dispersed (resulting in very low $PCF_{h}$), the system’s peakedness can remain low, even when user-level peakedness is high (as illustrated in Fig.3C-D). Transport planners and operators often rely on aggregated system-level departure time distribution to ascertain peak hours and demand loading patterns. However, this information alone does not offer insights into individual travel behaviours, such as users’ responses towards changes or disruptions in public transport services. Commuters with higher levels of departure time peakedness (represented as higher user-level $\psi_{h}$ values) would tend to have higher rigidity and lower flexibility in departure time choice and, thus, have higher rescheduling penalties and lower willingness to shift departure time. As such, service disruptions that require departure time changes can cause higher impact on users with stronger departure time peakedness than on those with lower peakedness levels. Assessing the disruption impacts based only on aggregated system-level departure distributions would not accurately capture the rescheduling penalties faced by individual users. In order to provide more user-centric transit services and improve user experience in managing disruptions, it is important to assess what a given disruption means to each user by considering individual users’ departure time peakedness characteristics.
To demonstrate how the proposed peakedness measurement framework can assist public transit operators in understanding the potential impact of disruptions on transit users, we estimate the three peakedness measures for each origin zone, defined as the 1-km radius circle centred around each bus stop. We identify origin zones with high peakedness in terms of each measure type, namely $\psi_{h}^{sys}$, $\overline{\psi_{h}}$, or $PCF_{h}$, within each of the three cities (Brisbane, Gold Coast, and Sunshine Coast). For a given measure type, we obtain the values of that peakedness measure for all origin zones within each city and determine the median value. Origin zones with peakedness measures greater than this median are considered to have high peakedness for that measure type. Fig.6 shows the spatial distribution of the origin zones with high peakedness in terms of $\psi_{h}^{sys}$, $\overline{\psi_{h}}$, and $PCF_{h}$ in the first, second, and third columns, respectively. Notably, there are considerable discrepancies between origin zones with high system-wide departure time peakedness ($\psi_{h}^{sys}$) and origin zones with high user-level departure time peakedness ($\overline{\psi_{h}}$). This observation empirically validates our conjecture that system-wide demand patterns do not consistently reflect their individual users’ departure time choice patterns. We argue that it is more informative to examine the two user-level components ($\overline{\psi_{h}}$ and $PCF_{h}$) separately, rather than relying on the aggregated measure ($\psi_{h}^{sys}$), as demonstrated in the fourth column of Fig.6. For instance, transit operators can prioritise origin zones with both high $\overline{\psi_{h}}$ and high $PCF_{h}$. Users in these zones exhibit higher departure time peakedness, leading to increased rescheduling penalties during service disruptions, and strong peak window coincidence, meaning more users will be affected if a service disruption occurs during the system-wide peak period. By prioritising these zones, transit operators can proactively address the challenges posed by disorders, minimising the overall impact on users and the whole system. In addition, transit operators may develop tailed strategies for different scenarios. For instance, in origin zones with high $\overline{\psi_{h}}$ and low $PCF_{h}$, users may exhibit high departure time peakedness individually, but their peak travel times are not as tightly clustered. Prioritising these zones might be beneficial in specific situations, such as disruptions that affect a broader time window. On the other hand, origin zones with low $\overline{\psi_{h}}$ and high $PCF_{h}$ could be prioritised when the goal is to minimise the impact on a large number of users within a specific time frame, even if their departure time distributions are more dispersed. Ultimately, the choice of prioritisation strategy among these scenarios depends on the nature of the service disruption and the specific performance goals that transit operators aim to achieve.
@@
Figure 6: The spatial distribution of origin zones with high peakedness in terms of three peakedness measures, namely $\psi_{h}^{sys}$, $\overline{\psi_{h}}$, and $PCF_{h}$.
Discussion

### v2

@@ 3.2 Measuring Departure Time Peakedness of Three Cities
Figure 7 shows these estimates for different peak window widths, $h\in\{5,10,20,30,45,60\}$. Across the cities, Brisbane exhibits overall higher $\overline{\psi_{h}}$ and higher $PCF_{h}$, and thus higher $\psi_{h}^{sys}$, than Gold Coast and Sunshine Coast in both the AM and PM periods. This is likely due to Brisbane’s higher population density and greater urbanisation, which create more structured, work-driven commuting patterns, resulting in more synchronised and concentrated departure times. On the other hand, the more leisure-oriented nature of Gold Coast and Sunshine Coast tends to result in more flexible and dispersed departure patterns. Between the AM and PM periods, peakedness measures are overall higher during the AM period than during the PM period in each city. This can be attributed to the stricter time constraints in the morning, with commuters needing to adhere to fixed start times for work or school, leading to a greater concentration of AM departure time distributions around their peaks. However, Brisbane’s $PCF_{h}$ measures display an interesting variation: $PCF_{h}$ of PM trips is higher than or similar to that of AM trips, indicating that users’ peak departure time windows coincide more strongly in the PM period than in the AM period, while the other two cities show much lower peak coincidence in the PM period. This can be explained by a more uniform end-of-day schedule across work sectors and after-school activities in Brisbane, resulting in greater synchronisation of departure times across individual users.
@@
3.3 Estimation of Optimal Peak Window
Figure 8 shows the estimation results of $\psi_{h}$ for different $h$ values, namely $h=5,10,\cdots,30$ minutes, for Brisbane and Gold Coast during two time-of-day periods (AM and PM). For each $h$, the distribution of the measured $\psi_{h}$ values across users is shown as a histogram, along with its mean ($\overline{\psi_{h}}$) and variance (Var). As expected, the distribution of the measured $\psi_{h}$ values shift to the right (i.e., the mean, $\overline{\psi_{h}}$, increases) as the window width, $h$, becomes larger. Meanwhile, the diversity or spread of the measured $\psi_{h}$ values (i.e., $\text{Var}[\psi_{h}]$) initially increases and then decreases as $h$ increases. Figure 9 provides a separate plot showing the change in $\text{Var}[\psi_{h}]$ with respect to $h$, where the $h$ value that maximises the variance (optimal peak window, $h^{*}$) is marked by a vertical dashed line.
The analysis indicates that the value of $h^{*}$ is consistent across the three study networks. Specifically, all three network indicate that $h^{*}$ is 15 minutes for the AM period and $h^{*}$ is 20-25 minutes for the PM period. The shorter optimal window for the AM period likely reflects more concentrated departure time distributions, as morning schedules are typically more constrained due to fixed start times for work or school. In contrast, departure times may be more spread out due to the varying end times of activities. When many individuals have tightly concentrated departure times (as in the AM period), the variance of $\psi_{h}$ values across users will peak at a shorter $h$ because smaller windows are sufficient to capture the high concentration of departures for most users. On the other hand, when the population has more dispersed or spread-out departure times (as in the PM period), the variance of $\psi_{h}$ across the population will reach its peak at a larger $h^{*}$ value because a larger $h$ window is needed to capture significant portions of the departure events across users. Despite the differences in the optimal window values between the AM and PM periods, the variance around those values is not significantly different. Overall, the analysis for Brisbane, Gold Coast, and Sunshine Coast reveals that the optimal window $h^{*}$ is consistently around 20 minutes, regardless of the time-of-day group or region. This suggests that there is a common temporal threshold where departure time peakedness is most distinguishable across the population. Thus, we consider using $h$=20 minutes as the most effective window for measuring trip concentrations ($\psi_{h}$), as it reveals the greatest diversity in departure time peakedness.
(a) Brisbane
(b) Gold Coast
Figure 8: Histograms of estimated $h$-min peak trip concentration ($\psi_{h}$) across users under different $h$ values ($h=5,10,\cdots,30$ minutes): the mean ($\overline{\psi_{h}}$) and variance (Var) of $\psi_{h}$ are presented in each histogram.
Figure 9: The variance ($\text{Var}[\psi_{h}]$) of estimated $h$-min peak concentration ($\psi_{h}$) across users under different $h$ values. The $h$ value that maximises the variance (i.e., optimal peak window, $h^{*}$) is marked by a vertical dashed line.
@@ 3.3 Estimation of Optimal Peak Window
Table 2 presents the descriptive statistics of the distribution of estimated $\psi_{h}$ values with $h$=20 minutes (i.e., $\psi_{20}$). Across the three cities, on average, approximately 61-63% of daily departure times occur during each user’s most concentrated 20-minute window in the AM period, and 46-49% in the PM period. Users in the top 5% for peakedness, represented by the 95th percentile, show that nearly 98-100% of their departures are concentrated within the peak 20-minute window in the AM and 92-97% in the PM. In contrast, users in the bottom 5% for peakedness, as indicated by the 5th percentile, have only 20-23% of their departure times within the peak 20-minute window for the AM period, and 16-17% for the PM period.
@@
3.4 Factors Affecting Departure Time Peakedness
@@
First, we explored how different passenger types shape departure time peakedness. Figure 10 compares the three peakedness measures across six passenger types: Adult, Tertiary Student, Pensioner, Secondary Student (15+ years), Child (5 to 14 years), and Senior. Notably, the child passenger type exhibits consistently high departure time peakedness across all three measures, particularly in user-level peakedness ($\overline{\psi_{h}}$). For example, in Brisbane, the $\overline{\psi_{h}}$ of children is nearly 0.8, indicating that most of their day-to-day departures occur within their most frequent 20-minute departure time window. This high concentration is likely due to the fixed start and end times of primary and secondary school, which children in the 5-14 year age group typically attend. Interestingly, the peak window coincidence ($PCF_{h}$) for the PM period is notably high for children in Brisbane ($PCF_{h}$=0.6) and Sunshine Coast ($PCF_{h}$=0.5), even surpassing the AM period. This could be attributed to the more uniform scheduling of after-school activities or programs across various areas in these cities, which may lead to a higher alignment of peak departure times among different users within the peak window. This finding provides some evidence to the role of deeply ingrained travel habits in shaping the travel choice of children (Mucelli et al., 2016) alongside other factors including service availability, perceived safety and accessibility.
Next, Figure 11 draws together a suite of situational variables known to impact bus ridership dynamics in the form of six scenarios: Weekday vs Weekend, Non School Holiday vs. School Holiday, and Rain vs. No Rain. We find that the departure time peakedness characteristics are relatively stable across these scenarios, with some subtle shifts. For example, in Brisbane and Gold Coast, weekend departure times are slightly less peaked than weekdays, likely due to more flexible weekend schedules with diverse non-work activities, leading to more dispersed travel times. In contrast, weekdays are dominated by structured work and school schedules, resulting in higher peakedness. Sunshine Coast does not show such distinction, possibly due to its tourism-driven, leisure-oriented travel patterns and flexible local employment, resulting in consistent travel behaviour across both weekdays and weekends.
@@
Figure 10: Comparison of peakedness measures ($h$ = 20 minutes) across different passenger types
Figure 11: Comparison of peakedness measures ($h$ = 20 minutes) under different scenarios
@@
3.5 Temporal Stability and Variability in Departure Time Peakedness: A Trend Analysis
Next we investigate long-term temporal dynamics of departure time peakedness to understand whether system- or user-level peakedness characteristics change over time. To explore this, we measured both $\psi_{h}^{sys}$ and $\psi_{h}$, with $h$ = 20 minutes, using data from a three-month moving window spanning over one year from July 2015 to June 2016. This results in a time-series with 10 data points, corresponding to the periods July-September 2015, August-October 2015, and so on, up to April-June 2016. These data points allow us to analyse trends or changes in the chosen peakedness measure over time.
To evaluate the presence of trends, we applied the Mann-Kendall (MK) Trend Test (Mann, 1945; Kendall, 1948), which assesses whether there is a statistically significant monotonic increase or decrease in the peakedness measures over time. As our time series data are relatively short, we used the original non-parametric MK test for analysing consistently increasing or decreasing trends (monotonic trends), which does not consider serial correlation or seasonality effects Hussain and Mahmud (2019). The MK test determines whether to reject the null hypothesis of no monotonic trend at a given significance level $\alpha$, which is set to 0.05 in this study.
For the system-level peakedness, the MK test results show that there is no monotonic trend in the series, which is true for all three cities for both AM and PM. Figure 12 shows the time-series of $\psi_{h}^{sys}$ used in the MK test, where the x-axis labels the first month of each three-month moving window in the format ‘YY/MM’ (Year/Month). The number of long-term regular users involved in constructing each system-level time-series ($n$) is shown in the parenthesis next to each city name within the legend. In all six cases, the value of $\psi_{h}^{sys}$ remains quite stable over the one-year period, indicating the absence of a significant trend in the system-level departure time peakedness.
In contrast, the MK test results reveal interesting trends in the user-level peakedness time-series data. Approximately 25% of the long-term regular users exhibit statistically significant trends in their $\psi_{h}$ time-series. Of these users, half show increasing trends and the other half show decreasing trends. This pattern is consistent across all three cities, regardless of the selected peak window width ($h$). The detailed results are displayed in Figure 13, which shows the proportion of user-OD pairs with decreasing or increasing $\psi_{h}$, both as a percentage of long-term regular users (indicated by taller bars in the background) and as a percentage of the original regular user base (indicated by shorter bars in the foreground). The results are obtained for the three cities using four different peak window widths ($h$ = 5, 10, 20, and 30 minutes). The first sub-figure shows the number of user-OD pairs exhibiting any monotonic trends (either decreasing or increasing $\psi_{h}$), while the second and third sub-figures show the numbers corresponding to increasing and decreasing trends, respectively.
@@ 3.5 Temporal Stability and Variability in Departure Time Peakedness: A Trend Analysis
To sum, at the system level, our analysis results show no significant monotonic trends in any city (Brisbane, Gold Coast, or Sunshine Coast), indicating that system-level departure time peakedness remains stable over time. This suggests that the collective travel behaviour of long-term regular users does not exhibit major fluctuations in departure time patterns throughout the year. However, at the user level, the analysis reveals notable dynamics. Approximately 25% of long-term regular users (those making trips consistently throughout the year) exhibit significant trends in their $\psi_{h}$ time-series. Interestingly, these trends are balanced: about half of the users show increasing peakedness (with more concentrated and regular departure times over time), while the other half display decreasing peakedness (with more dispersed and flexible travel patterns over time). Some users exhibit gradual changes, while others experience sudden shifts. Overall, this trend analysis highlights the stability of system-level peakedness, while revealing the evolving nature of individual-level departure time behaviour. It also suggests that habitual travel patterns, though resistant to short-term situational factors like weather or holidays (as shown in Section 3.4), may shift over the long term due to larger-scale factors such as seasonality, service changes, or shifts in user characteristics.
3.6 Applications of Peakedness Measures
One of the most critical applications of our peakedness measurement framework is in disruption management. Commuters with higher levels of departure time peakedness (represented as higher user-level $\psi_{h}$ values) would tend to have higher rigidity and lower flexibility in departure time choice and, thus, have higher rescheduling penalties and lower willingness to shift departure time. As such, service disruptions that require departure time changes can cause higher impact on users with stronger departure time peakedness than on those with lower peakedness levels. Traditional transit planning and operations often rely on system-level departure time distributions to identify peak hours and assess demand patterns. However, these aggregated patterns do not necessarily reflect individual user behaviours, particularly their flexibility or rigidity in departure time choices.
@@ 3.6 Applications of Peakedness Measures
Our method reveals that user-level peakedness and system-level peakedness do not always exhibit a direct correlation. In some cases, areas with users whose departure time distributions have a high degree of peakedness may still result in a combined departure time distribution at the system level that shows low peakedness. This discrepancy arises because the system-level distribution peakedness ($\psi_{h}^{sys}$) is influenced by both the peakedness of individual departure time distributions ($\overline{\psi_{h}}$) and the alignment of their peak time windows ($PCF_{h}$). In situations where the peak windows of individuals are widely dispersed (resulting in very low $PCF_{h}$), the system’s peakedness can remain low, even when user-level peakedness is high (as illustrated in Figure 3(c)-(d). Assessing the disruption impacts based only on aggregated system-level departure distributions would not accurately capture the rescheduling penalties faced by individual users. In order to provide more user-centric transit services and improve user experience in managing disruptions, it is important to assess what a given disruption means to each user by considering individual users’ departure time peakedness characteristics.
@@
Figure 16: The spatial distribution of origin zones with high departure-time peakedness in terms of three peakedness measures, namely $\psi_{h}^{sys}$, $\overline{\psi_{h}}$, and $PCF_{h}$ (AM period).
To demonstrate how the proposed peakedness measurement framework can assist public transit operators in understanding the potential impact of disruptions on transit users, we estimate the three peakedness measures for each origin zone, defined as the 1-km radius circle centred around each bus stop. We identify origin zones with high peakedness in terms of each measure type, namely $\psi_{h}^{sys}$, $\overline{\psi_{h}}$, or $PCF_{h}$, within each of the three cities (Brisbane, Gold Coast, and Sunshine Coast) during the AM period. For a given measure type, we obtain the values of that peakedness measure for all origin zones within each city and determine the median value. Origin zones with peakedness measures greater than this median are considered to have high peakedness for that measure type. Figure 16 shows the spatial distribution of the origin zones with high peakedness in terms of $\psi_{h}^{sys}$, $\overline{\psi_{h}}$, and $PCF_{h}$ in the first, second, and third rows, respectively. Notably, there are considerable discrepancies between origin zones with high system-wide departure time peakedness ($\psi_{h}^{sys}$) and origin zones with high user-level departure time peakedness ($\overline{\psi_{h}}$). This observation empirically validates our conjecture that system-wide demand patterns do not consistently reflect their individual users’ departure time choice patterns. We argue that it is more informative to examine the two user-level components ($\overline{\psi_{h}}$ and $PCF_{h}$) separately, rather than relying on the aggregated measure ($\psi_{h}^{sys}$), as demonstrated in the fourth row of Figure 16. For instance, transit operators can prioritise origin zones with both high $\overline{\psi_{h}}$ and high $PCF_{h}$. Users in these zones exhibit higher departure time peakedness, leading to increased rescheduling penalties during service disruptions, and strong peak window coincidence, meaning more users will be affected if a service disruption occurs during the system-wide peak period. By prioritising these zones, transit operators can proactively address the challenges posed by disorders, minimising the overall impact on users and the whole system. In addition, transit operators may develop tailored strategies for different scenarios. For instance, in origin zones with high $\overline{\psi_{h}}$ and low $PCF_{h}$, users may exhibit high departure time peakedness individually, but their peak travel times are not as tightly clustered. Prioritising these zones might be beneficial in specific situations, such as disruptions that affect a broader time window. On the other hand, origin zones with low $\overline{\psi_{h}}$ and high $PCF_{h}$ could be prioritised when the goal is to minimise the impact on a large number of users within a specific time frame, even if their departure time distributions are more dispersed. Ultimately, the choice of prioritisation strategy among these scenarios depends on the nature of the service disruption and the specific performance goals that transit operators aim to achieve.
We have also conducted an analysis investigating destination zones to provide a more comprehensive understanding. This additional analysis, presented in F, examines arrival time peakedness and reveals similar discrepancies between system-wide and user-level peakedness measures as observed in the origin-based analysis. By considering both origin and destination zones, transit operators can identify areas with high peakedness in both departure and arrival patterns, allowing them to develop more tailored disruption management strategies that account for rigid travel behaviours at both ends of the journey.
4 Conclusion